OpenStudy (anonymous):

For the simple harmonic motion equation d= 5 sin(pi/4 t) , what is the period

OpenStudy (anonymous):

w = pi /4 = 2 pi f 1 /4 = 2 f f = 1/8 = angular frequency = 1 / period period = 8 unit

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Anytime :)